OpenStudy (anonymous):

The probability of winning a game is 15%. If you play 20 times, how many times should you expect to win A.5 times B.3 times C.6 times D.15 times

OpenStudy (welshfella):

20 * 15 ------ 100

OpenStudy (anonymous):

what does that mean?

OpenStudy (dahlindubs11):

no... its .15*20=3

OpenStudy (misty1212):

hmm seems unlikely to me

OpenStudy (misty1212):

\(15\%\) of \(20\) is \(3\)

OpenStudy (welshfella):

20 * 15 = 300 - now divide by 100

OpenStudy (anonymous):

@thanks